OpenStudy (anonymous):

-3/2 k + -5/8 = -8--9/7 k plz

OpenStudy (anonymous):

You want to isolate k. So put k on one side and the numbers without k on the other side.

OpenStudy (anonymous):

Linear Equations

OpenStudy (anonymous):

How in the world did you type that so quickly?

OpenStudy (anonymous):

You typed that in one minute?

OpenStudy (anonymous):

it wrong

OpenStudy (anonymous):

i put the answer and it wrong

OpenStudy (anonymous):

I got -22.25

OpenStudy (anonymous):

that the question

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

(-3)/(2)*k+(-5)/(3)=-8-(-9)/(7)*k Remove all extra parentheses from the expression. -(3)/(2)*k+(-5)/(3)=-8-(-9)/(7)*k Remove all extra parentheses from the expression. -(3)/(2)*k-(5)/(3)=-8-(-9)/(7)*k Multiply -(3)/(2) by k to get -(3k)/(2). -(3k)/(2)-(5)/(3)=-8-(-9)/(7)*k Remove all extra parentheses from the expression. -(3k)/(2)-(5)/(3)=-8+(9)/(7)*k Multiply (9)/(7) by k to get (9k)/(7). -(3k)/(2)-(5)/(3)=-8+(9k)/(7) Since -(5)/(3) does not contain the variable to solve for, move it to the right-hand side of the equation by adding (5)/(3) to both sides. -(3k)/(2)=(5)/(3)+(9k)/(7)-8 Simplify the right-hand side of the equation. -(3k)/(2)=(9k)/(7)-(19)/(3) Find the LCD (least common denominator) of -(3k)/(2)+(9k)/(7)-(19)/(3). Least common denominator: 42 Multiply each term in the equation by 42 in order to remove all the denominators from the equation. -(3k)/(2)*42=(9k)/(7)*42-(19)/(3)*42 Simplify the left-hand side of the equation by canceling the common factors. -63k=(9k)/(7)*42-(19)/(3)*42 Simplify the right-hand side of the equation by simplifying each term. -63k=54k-266 Since 54k contains the variable to solve for, move it to the left-hand side of the equation by subtracting 54k from both sides. -63k-54k=-266 Since -63k and -54k are like terms, subtract 54k from -63k to get -117k. -117k=-266 Divide each term in the equation by -117. -(117k)/(-117)=-(266)/(-117) Simplify the left-hand side of the equation by canceling the common factors. k=-(266)/(-117) Simplify the right-hand side of the equation by simplifying each term. k=(266)/(117)

OpenStudy (anonymous):

yea thank u

OpenStudy (anonymous):

what about this one

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

(1)/(2)*x+(1)/(3)=-1((5)/(6)*x+3) Multiply 1 by x to get x. (x)/(2)+(1)/(3)=-1((5)/(6)*x+3) Multiply 5 by x to get 5x. (x)/(2)+(1)/(3)=-1((5x)/(6)+3) Multiply -1 by each term inside the parentheses ((5x)/(6)+3). (x)/(2)+(1)/(3)=-1((5x)/(6))-1(3) Multiply -1 by the (5x)/(6) inside the parentheses. (x)/(2)+(1)/(3)=-1*(5x)/(6)-1(3) Multiply -1 by 5x to get -5x. (x)/(2)+(1)/(3)=-(5x)/(6)-1(3) Multiply -1 by the 3 inside the parentheses. (x)/(2)+(1)/(3)=-(5x)/(6)-1*3 Multiply -1 by 3 to get -3. (x)/(2)+(1)/(3)=-(5x)/(6)-3 Since (1)/(3) does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting (1)/(3) from both sides. (x)/(2)=-(1)/(3)-(5x)/(6)-3 Reorder the polynomial -(1)/(3)-(5x)/(6)-3 alphabetically from left to right, starting with the highest order term. (x)/(2)=-(5x)/(6)-(1)/(3)-3 To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 3. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (x)/(2)=-(5x)/(6)-3*(3)/(3)-(1)/(3) Multiply -3 by 3 to get -9. (x)/(2)=-(5x)/(6)-(9)/(3)-(1)/(3) Combine the numerators of all fractions that have common denominators. (x)/(2)=-(5x)/(6)+(-9-1)/(3) Subtract 1 from -9 to get -10. (x)/(2)=-(5x)/(6)+(-10)/(3) Move the minus sign from the numerator to the front of the expression. (x)/(2)=-(5x)/(6)-(10)/(3) Find the LCD (least common denominator) of (x)/(2)-(5x)/(6)-(10)/(3). Least common denominator: 6 Multiply each term in the equation by 6 in order to remove all the denominators from the equation. (x)/(2)*6=-(5x)/(6)*6-(10)/(3)*6 Cancel the common factor of 2 in the denominator of the first term (x)/(2) and the second term 6. (x)/(<X>2<x>)*<X>6<x>^(3)=-(5x)/(6)*6-(10)/(3)*6 Reduce the expression by removing the common factor of 2 in the denominator of the first term (x)/(2) and the second term 6. x*3=-(5x)/(6)*6-(10)/(3)*6 Multiply x by 3 to get 3x. 3x=-(5x)/(6)*6-(10)/(3)*6 Cancel the common factor of 6 in the denominator of the first term -(5x)/(6) and the second term 6. 3x=-(5x)/(<X>6<x>)*<X>6<x>-(10)/(3)*6 Reduce the expression by removing the common factor of 6 in the denominator of the first term -(5x)/(6) and the second term 6. 3x=-5x*1-(10)/(3)*6 Cancel the common factor of 3 in the denominator of the first term -(10)/(3) and the second term 6. 3x=-5x*1-(10)/(<X>3<x>)*<X>6<x>^(2) Reduce the expression by removing the common factor of 3 in the denominator of the first term -(10)/(3) and the second term 6. 3x=-5x*1-10*2 Multiply -5x by 1 to get -5x. 3x=-5x-10*2 Multiply -10 by 2 to get -20. 3x=-5x-20 Since -5x contains the variable to solve for, move it to the left-hand side of the equation by adding 5x to both sides. 3x+5x=-20 According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, x is a factor of both 3x and 5x. (3+5)x=-20 Add 5 to 3 to get 8. (8)x=-20 Remove the parentheses. 8x=-20 Divide each term in the equation by 8. (8x)/(8)=-(20)/(8) Cancel the common factor of 8 in (8x)/(8). (<X>8<x>x)/(<X>8<x>)=-(20)/(8) Remove the common factors that were cancelled out. x=-(20)/(8) Cancel the common factor of 4 in -(20)/(8) since -(20)/(8)=((-5*4))/((2*4)). x=-(^(5)<X>20<x>)/(2<X>8<x>) Remove the common factors that were cancelled out. x=-(5)/(2)

OpenStudy (anonymous):

how u do that fast??

OpenStudy (anonymous):

How was that fast? lol

OpenStudy (anonymous):

well thanks for helping but i got the last one sorry i mean u type fast to write all that..

OpenStudy (anonymous):

You should try it this time :)

OpenStudy (anonymous):

okay i well try but i dont get how to move the fraction to the other side that the part i miss up

OpenStudy (anonymous):

i got 1/4

OpenStudy (anonymous):

???

OpenStudy (anonymous):

(1)/(3)*x+(1)/(2)=-3((3)/(4)*x-1) Multiply 1 by x to get x. (x)/(3)+(1)/(2)=-3((3)/(4)*x-1) Multiply 3 by x to get 3x. (x)/(3)+(1)/(2)=-3((3x)/(4)-1) Multiply -3 by each term inside the parentheses ((3x)/(4)-1). (x)/(3)+(1)/(2)=-3((3x)/(4))-3(-1) Multiply -3 by the (3x)/(4) inside the parentheses. (x)/(3)+(1)/(2)=-3*(3x)/(4)-3(-1) Multiply -3 by 3x to get -9x. (x)/(3)+(1)/(2)=-(9x)/(4)-3(-1) Multiply -3 by the -1 inside the parentheses. (x)/(3)+(1)/(2)=-(9x)/(4)-3*-1 Multiply -3 by -1 to get 3. (x)/(3)+(1)/(2)=-(9x)/(4)+3 Since (1)/(2) does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting (1)/(2) from both sides. (x)/(3)=-(1)/(2)-(9x)/(4)+3 Reorder the polynomial -(1)/(2)-(9x)/(4)+3 alphabetically from left to right, starting with the highest order term. (x)/(3)=-(9x)/(4)-(1)/(2)+3 To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 2. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (x)/(3)=-(9x)/(4)+3*(2)/(2)-(1)/(2) Multiply 3 by 2 to get 6. (x)/(3)=-(9x)/(4)+(6)/(2)-(1)/(2) Combine the numerators of all fractions that have common denominators. (x)/(3)=-(9x)/(4)+(6-1)/(2) Subtract 1 from 6 to get 5. (x)/(3)=-(9x)/(4)+(5)/(2) Find the LCD (least common denominator) of (x)/(3)-(9x)/(4)+(5)/(2). Least common denominator: 12 Multiply each term in the equation by 12 in order to remove all the denominators from the equation. (x)/(3)*12=-(9x)/(4)*12+(5)/(2)*12 Cancel the common factor of 3 in the denominator of the first term (x)/(3) and the second term 12. (x)/(<X>3<x>)*<X>12<x>^(4)=-(9x)/(4)*12+(5)/(2)*12 Reduce the expression by removing the common factor of 3 in the denominator of the first term (x)/(3) and the second term 12. x*4=-(9x)/(4)*12+(5)/(2)*12 Multiply x by 4 to get 4x. 4x=-(9x)/(4)*12+(5)/(2)*12 Cancel the common factor of 4 in the denominator of the first term -(9x)/(4) and the second term 12. 4x=-(9x)/(<X>4<x>)*(<X>12<x>^(3)5)/(2)*12 Reduce the expression by removing the common factor of 4 in the denominator of the first term -(9x)/(4) and the second term 12. 4x=-9x*3+(5)/(2)*12 Cancel the common factor of 2 in the denominator of the first term (5)/(2) and the second term 12. 4x=-9x*3+(5)/(<X>2<x>)*<X>12<x>^(6) Reduce the expression by removing the common factor of 2 in the denominator of the first term (5)/(2) and the second term 12. 4x=-9x*3+5*6 Multiply -9x by 3 to get -27x. 4x=-27x+5*6 Multiply 5 by 6 to get 30. 4x=-27x+30 Since -27x contains the variable to solve for, move it to the left-hand side of the equation by adding 27x to both sides. 4x+27x=30 According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, x is a factor of both 4x and 27x. (4+27)x=30 Add 27 to 4 to get 31. (31)x=30 Remove the parentheses. 31x=30 Divide each term in the equation by 31. (31x)/(31)=(30)/(31) Cancel the common factor of 31 in (31x)/(31). (<X>31<x>x)/(<X>31<x>)=(30)/(31) Remove the common factors that were cancelled out. x=30/31